Sushil is now 4times as old as sunil.Five years ago sushil was 7 times as old as sunil was then.Find the present age of each of them .this is how I did- when I tried I got-
4+x= 5 (7+x)
= 4+x= 35+5x
= 35-4=5x-x
31=4x
X= 4/31



Answer by Cromlix(4381)

Hi there,
Sushil is now 4times as old as sunil.
Sush = 4Sun
Five years ago sushil was 7 times as
old as sunil was then.
Sush - 5 = 7(Sun - 5)
Sush - 5 = 7Sun - 35
Replace Sush with 4Sun from top equation
4Sun - 5 = 7Sun - 35
Collect like terms
4Sun - 7Sun = -35 + 5
-3Sun = -30
Multiply both sides by -1
3Sun = 30
Sun = 10
Sunhil is currently 10 years old
Sushil is currently 40 years old.
Hope this helps :-)